# Compositional Generative Multiphysics and Multi-component Simulation

### v1

@@ Abstract
Multiphysics simulation, which models the interactions between multiple physical processes, and multi-component simulation of complex structures are critical in fields like nuclear and aerospace engineering. Previous studies often rely on numerical solvers or machine learning-based surrogate models to solve or accelerate these simulations. However, multiphysics simulations typically require integrating multiple specialized solvers—each responsible for evolving a specific physical process—into a coupled program, which introduces significant development challenges. Furthermore, no universal algorithm exists for multi-component simulations, which adds to the complexity. Here we propose compositional Multiphysics and Multi-component Simulation with Diffusion models (MultiSimDiff) to overcome these challenges. During diffusion-based training, MultiSimDiff learns energy functions modeling the conditional probability of one physical process/component conditioned on other processes/components. In inference, MultiSimDiff generates coupled multiphysics solutions and multi-component structures by sampling from the joint probability distribution, achieved by composing the learned energy functions in a structured way. We test our method in three tasks. In the reaction-diffusion and nuclear thermal coupling problems, MultiSimDiff successfully predicts the coupling solution using decoupled data, while the surrogate model fails in the more complex second problem. For the thermal and mechanical analysis of the prismatic fuel element, MultiSimDiff trained for single component prediction accurately predicts a larger structure with 64 components, reducing the relative error by 40.3% compared to the surrogate model.
@@ 1 INTRODUCTION
Multiphysics simulation involves the concurrent modeling of multiple physical processes—such as heat conduction, fluid flow, and structural mechanics—within a single simulation framework to accurately capture the coupling effects between different physical processes. Similarly, multi-component simulation focuses on simulating complex structures composed of multiple similar components. Component is defined as: a repeatable basic unit that makes up a complete structure. For example, the reactor core typically consists of hundreds or thousands of fuel elements arranged in a square or hexagonal pattern. These simulations are essential across various scientific and engineering disciplines, including nuclear engineering (Ma et al., 2022; Chen et al., 2021), aerospace engineering (Candeo et al., 2011; Wang et al., 2023a), civil engineering (Sun et al., 2017; Meyer et al., 2022), and automotive industry (Ragone et al., 2021). Despite their significance, both multiphysics and multi-component simulations share a common challenge: while simulating individual components or physical processes is relatively straightforward, modeling the entire system with all its interactions is vastly more complex.
@@
Numerous numerical algorithms have been developed for multiphysics simulation, which are broadly categorized into loose coupling and tight coupling (Hales et al., 2015). Loose coupling involves solving each physical process independently while treating the others as constant. Solutions for one physical process are iteratively transferred to related physical processes until convergence is achieved, often using techniques like operator splitting (MacNamara & Strang, 2016) and Picard iteration (Terlizzi & Kotlyar, 2022). Tight coupling, on the other hand, assembles equations of all physical processes into a large system, solving them simultaneously (Knoll & Keyes, 2004). While this method can potentially yield more accurate results, it encounters challenges such as high computational costs, varying spatial and temporal resolutions, and differing numerical methods across physical processes, leading to a more common use of loose coupling in engineering applications. In multi-component simulation, directly simulating the overall structure requires high computational cost and may encounter difficulties in convergence due to the increase in degrees of freedom. Substructure methods have been used in fields like nuclear engineering (Chen et al., 2021) and civil engineering (Sun et al., 2017) to reduce modeling and computational costs for repetitive components.
@@
Despite advances in numerical algorithms, several significant challenges remain. In multiphysics simulations, considerable time and effort are required to develop programs that couple different specialized solvers. Furthermore, the complexity of the system increases due to coupling, requiring more computing resources. While some studies employ machine learning-based surrogate models to accelerate multiphysics simulations (Sobes et al., 2021; Park et al., 2021), these models still depend on coupled data for training, which necessitates the prior development of coupled numerical solution programs. In the case of multi-component simulations, the substructure method has primarily been applied to mechanical problems, with no widely applicable general method for multi-component systems. Consequently, current approaches often rely on selecting representative units or implementing simplifications for the analysis of complex structures, which can compromise the accuracy and scope of the simulation.
@@
Figure 1: MultiSimDiff schematic. Our proposed algorithm can use models trained with decoupled data to predict coupled solutions (top) and use models trained with small structure simulation data to predict solutions for large structures (here 64 components) (bottom).
@@
To address these challenges, we propose compositional Multiphysics and Multi-component Simulation with Diffusion models (MultiSimDiff). The core innovation of MultiSimDiff is its treatment of multiphysics and multi-component simulations as generative probabilistic modeling, where interactions between multiple physical processes or components are captured through composing learned energy functions conditioned on others in a structured way. In multiphysics simulation, MultiSimDiff generates coupled solutions (accounting for interactions between different physical processes) from decoupled data (assuming other fields are known and focused on solving a single field) by modeling the solutions of physical processes as a joint probability distribution. The solution for each individual process is treated as a conditional probability distribution, based on Bayes’ theorem. By training diffusion models (Ho et al., 2020) on decoupled data, we capture these conditional distributions. During inference, the model combines these distributions and performs reverse diffusion to produce the coupled solution. For multi-component simulations, MultiSimDiff models each component’s solution as a conditional probability distribution using the local Markov property, conditioned on neighboring components. By training diffusion models on small structures, we create conditional models for individual components. During inference, reverse diffusion is applied iteratively across all components, yielding the solution for the entire structure. We have mathematically derived the principles why MultiSimDiff can obtain coupled solutions and large structure solutions in Sections 3.1 and 3.2. A schematic of MultiSimDiff is provided in Fig. 1.
@@
We illustrate the promise of this approach through three challenging tasks. First, we demonstrate its capability for multiphysics simulation by applying it to coupled reaction-diffusion equations and nuclear thermal coupling combined with conjugate heat transfer. Second, we verify its capability in multi-component simulation through thermal and mechanical analysis of prismatic fuel elements.
@@
Concretely, our contributions are threefold: (1) We introduce a novel approach, MultiSimDiff, for multiphysics and multi-component simulations, framing the problem in terms of joint probabilistic modeling. By training on decoupled (small structure) training data, MultiSimDiff can generate coupled (large structure) solutions. (2) We create and open-source benchmark datasets for both multiphysics and multi-component simulations, providing a valuable resource for future research. (3) Our method demonstrates success in both domains. For multiphysics simulation, MultiSimDiff accurately predicts coupled solutions in complex problems where surrogate models fail. In multi-component simulations, MultiSimDiff, trained on single components, accurately predicts larger structures with up to 64 components, reducing relative error by 40.3% compared to surrogate models.
@@ 2 RELATED WORK
Multi-component simulation. To our knowledge, there do not exist utilized machine learning methods specifically designed for multi-component simulation. A relevant study is the CoAE-MLSim algorithm (Ranade et al., 2021). This algorithm combines neural networks with numerical iteration. It first partitions the computational domain into multiple subdomains, and then trains a neural network to learn the flux conversation between subdomains. During inference, the neural network with flux conservation is applied sequentially in each subdomain, looping until convergence. We further extend this algorithm to multi-component simulation and use it as a baseline. Besides, graph neural network (GNN) (Wu et al., 2020) can learn on small graphs and inference on larger graphs (Xu et al., 2019); Graph Transformer (Kreuzer et al., 2021) employs the Laplacian matrix of a graph to characterize its structure, and by leveraging the Transformer architecture, it achieves learning on graphs. We also compare MultiSimDiff with GNN and Graph Transformer.
Compositional models. Recent research has extensively explored the compositional combination of generative models for various applications, including 2D image synthesis (Du et al., 2020; Liu et al., 2021; Nie et al., 2021; Liu et al., 2022; Wu et al., 2022; Du et al., 2023; Wang et al., 2023c), 3D synthesis (Po & Wetzstein, 2024), video synthesis (Yang et al., 2023a), multimodal perception (Li et al., 2022), trajectory planning (Du et al., 2019; Urain et al., 2023; Gkanatsios et al., 2024; Yang et al., 2023b), inverse design (Wu et al., 2024b), and hierarchical decision making (Ajay et al., 2024). A particularly effective approach for combining predictive distributions from local experts is the product of experts framework (Hinton, 2002; Cohen et al., 2020; Kant et al., 2024; Tautvaišas & Žilinskas, 2023). Their focus is on how a single object is influenced by multiple factors, such as generating images that meet various requirements in image generation (Du et al., 2023) or enhancing the lift-to-drag ratio under the influence of two wings in inverse design (Wu et al., 2024b). However, our problem involves multiple objects, such as multiple physical processes and components, requiring the capture of interactions between these fields or components. Existing research is not applicable to multiphysics and multi-component simulation. To the best of our knowledge, we are the first to introduce a compositional generative approach to multiphysics and multi-component simulations, demonstrating how this framework enables generalization to far more complex simulation tasks than those encountered during training.
@@ 3 METHOD
In this section, we introduce the principle of MultiSimDiff solving multiphysics and multi-component simulation in section 3.1 and section 3.2, respectively.
@@ 3.1 MULTIPHYSICS SIMULATION
Consider a complex multiphysics simulation problem that consists of multiple physical processes : $z=(z_{1},z_{2},\ldots,z_{n})$, where each $z_{i}$ may contain one or more fields. For example, the mechanics contains the stress and strain fields in three directions. Each process $z_{i}$ has its own governing equation which depends on other processes , and solving equations for other processes also requires that process. Therefore, all equations must be solved simultaneously to achieve the most accurate representation of the physical system.
@@
$E(z)+\mathrm{log}Z=[E(z_{i}\mid z_{\neq i})+\mathrm{log}Z(z_{\neq i})]+[E(z_{\neq i})+\mathrm{log}Z_{\neq i}]$ (5)
@@
$L_{\text{MSE}}=||\epsilon-\epsilon_{\theta}(\sqrt{1-\beta_{s}}z_{i}+\sqrt{\beta}_{s}\epsilon;z_{\neq i},s)||_{2}^{2},\,\,\epsilon\sim\mathcal{N}(0,I)$ (7)
where the denoising network $\epsilon_{\theta}(\cdot)$ corresponds to the gradient of the energy function $\nabla_{z}E_{\theta}(\cdot)$ (Du et al., 2023). During inference, we sample from the joint distribution $p(z_{1},z_{2},...z_{n})$ via (Ho et al., 2020):
$z_{i,s-1}=\frac{1}{\sqrt{\alpha_{s}}}\left(z_{i,s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}^{i}(z_{i,s}\mid z_{\neq i}^{e},s)\right)+\sigma_{s}w,\,\,\,w\sim\mathcal{N}(0,I)$ (8)
@@
for $s=S,S-1,...1$ and $i=1,2,...n$. Here, ${z}_{i}^{e}$ represents the estimated value for the $i$th field $z_{i}$, $z^{e}_{\neq i}=(z^{e}_{1},...z^{e}_{i-1},z^{e}_{i-1},...z^{e}_{n})$, and $\sigma_{s}$ is the noise level.
@@
Algorithm 1 Algorithm for multiphysics simulation by MultiSimDiff.
@@
1:  $z_{i}^{e}\sim\mathcal{N}(0,I)$  // initialize estimated fields $z_{i}^{e}$// add an external loop to improve the estimated fields $z_{i}^{e}$:
@@
3:     $\hat{z}_{i}^{e}\leftarrow z_{i}^{e}$  // update previous estimated fields $\hat{z}_{i}^{e}$
4:     $z_{i}^{e}\sim\mathcal{N}(0,\bf I)$  // initialize current estimated fields $z_{i}^{e}$
5:     $z_{i,S}\sim\mathcal{N}(0,\bf I)$  // initialize physical fields $z_{i}$// denoising cycle of diffusion model:
6:     for $s=S,...,1$ do
7:        $\lambda=1-\frac{s}{S}$ if $k>1$ else 1  // define the weights of $\hat{z}_{i}^{e}$ and $z_{i}^{e}$// loops for each physical process:
8:        for $i=1,...,N$ do
9:           $w\sim\mathcal{N}(0,\bf I)$// use weighted estimated fields as conditions for single step denoising:
10:           $z_{i,s-1}=\frac{1}{\sqrt{\alpha_{s}}}(z_{i,s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}^{i}(z_{i,s}\mid\lambda z_{\neq i}^{e}+(1-\lambda)\hat{z}_{\neq i}^{e},C,s))+\sigma_{s}w$// update the estimation of current field:
11:           $z_{i}^{e}=\frac{1}{\sqrt{\overline{\alpha}_{s}}}(z_{i,s}-\sqrt{1-\overline{\alpha}_{s}}\epsilon_{\theta}^{i}(z_{i,s}\mid\lambda z_{\neq i}^{e}+(1-\lambda)\hat{z}_{\neq i}^{e},C,s))$
12:        end for

### v2

# M2PDE: Compositional Generative Multiphysics and Multi-component PDE Simulation

@@ Abstract
Multiphysics simulation, which models the interactions between multiple physical processes, and multi-component simulation of complex structures are critical in fields like nuclear and aerospace engineering. Previous studies use numerical solvers or ML-based surrogate models for these simulations. However, multiphysics simulations typically require integrating multiple specialized solvers—each for a specific physical process—into a coupled program, which introduces significant development challenges. Furthermore, existing numerical algorithms struggle with highly complex large-scale structures in multi-component simulations. Here we propose compositional Multiphysics and Multi-component PDE Simulation with Diffusion models (M2PDE) to overcome these challenges. During diffusion-based training, M2PDE learns energy functions modeling the conditional probability of one physical process/component conditioned on other processes/components. In inference, M2PDE generates coupled multiphysics and multi-component solutions by sampling from the joint probability distribution. We evaluate M2PDE on two multiphysics tasks—reaction-diffusion and nuclear thermal coupling—where it achieves more accurate predictions than surrogate models in challenging scenarios. We then apply it to a multi-component prismatic fuel element problem, demonstrating that M2PDE scales from single-component training to a 64-component structure and outperforms existing domain-decomposition and graph-based approaches. The code is at github.com/AI4Science-WestlakeU/M2PDE.
Machine Learning, ICML
@@ 1 Introduction
Multiphysics simulation involves the concurrent modeling of multiple physical processes—such as heat conduction, fluid flow, and structural mechanics—within a single simulation framework to accurately capture the coupling effects between different physical processes. Similarly, multi-component simulation focuses on simulating complex structures composed of multiple similar components. Component is defined as: a repeatable basic unit that makes up a complete structure. For example, the reactor core typically consists of hundreds or thousands of fuel elements arranged in a square or hexagonal pattern. These simulations are essential across various scientific and engineering disciplines, including nuclear engineering (Ma et al., 2022; Chen et al., 2021), aerospace engineering (Candeo et al., 2011; Wang et al., 2023a), civil engineering (Sun et al., 2017; Meyer et al., 2022), and automotive industry (Ragone et al., 2021). Despite their significance, both multiphysics and multi-component partial differential equation (PDE) simulations share a common challenge: while simulating individual components or physical processes is relatively straightforward, modeling the entire system with all its interactions is vastly more complex.
@@
Numerous numerical algorithms have been developed for multiphysics simulation, which are broadly categorized into loose coupling and tight coupling (Hales et al., 2015). Loose coupling solves each physical process independently, iteratively transferring solutions until convergence, using methods like operator splitting (MacNamara & Strang, 2016) and Picard iteration (Terlizzi & Kotlyar, 2022). Tight coupling solves all processes simultaneously in a large system (Knoll & Keyes, 2004), which can be more accurate but faces challenges such as high computational costs, varying spatial and temporal resolutions, and differing numerical methods across physical processes. Thus, loose coupling is more commonly used in engineering applications. In multi-component simulation, directly simulating the overall structure requires high computational cost and may encounter difficulties in convergence due to the increase in degrees of freedom. Substructure methods have been used in fields like nuclear engineering (Chen et al., 2021) and civil engineering (Sun et al., 2017) to reduce modeling and computational costs for repetitive components.
@@
Despite advances in numerical algorithms, several significant challenges remain in engineering applications. In multiphysics simulations, considerable time and effort are required to develop programs that couple different specialized solvers. Furthermore, the complexity of the system increases due to coupling, requiring more computing resources. While some studies employ machine learning-based surrogate models to accelerate simulations (Sobes et al., 2021; Park et al., 2021), these models still depend on coupled data for training, which necessitates the prior development of coupled numerical solution programs. In the case of multi-component simulations, the substructure method has primarily been applied to mechanical problems, with no widely applicable general method for multi-component systems. Consequently, current methods often rely on selecting representative units for detailed analysis or introducing simplifications to the overall structure, which may compromise the accuracy and scope of the simulations.
@@
Figure 1: M2PDE schematic. Our proposed algorithm can use models trained with decoupled data to predict coupled solutions (top) and use models trained with small structure to predict large structures (here 64 components) (bottom).
@@
To address these challenges in engineering applications, we propose compositional Multiphysics and Multi-component PDE Simulation with Diffusion models (M2PDE). The core innovation of M2PDE is its treatment of multiphysics and multi-component PDE simulations as generative probabilistic modeling, where interactions between multiple physical processes or components are captured through composing learned energy functions conditioned on others in a structured way. In multiphysics simulation, M2PDE generates coupled solutions (accounting for interactions between different physical processes) from decoupled data (assuming other fields are known and focused on solving a single field) by modeling the solutions of physical processes as a joint probability distribution. The solution for each individual process is treated as a conditional probability distribution, based on Bayes’ theorem. By training diffusion models (Ho et al., 2020) on decoupled data, we capture these conditional distributions. During inference, the model combines these distributions and performs reverse diffusion to produce the coupled solution. For multi-component simulations, M2PDE models each component’s solution as a conditional probability distribution using the local Markov property, conditioned on neighboring components. By training diffusion models on small structures, we create conditional models for individual components. During inference, reverse diffusion is applied iteratively across all components, yielding the solution for the entire structure. We have mathematically derived the principles why M2PDE can obtain coupled solutions and large structure solutions in Sections 3.1 and 3.2. A schematic of M2PDE is provided in Fig. 1.
@@
We illustrate the promise of this approach through simplified problems in practical engineering, both of which are important in nuclear engineering and has widespread implications across engineering. First, we demonstrate its capability for multiphysics simulation by applying it to coupled reaction-diffusion equations and nuclear thermal coupling combined with conjugate heat transfer. Second, we verify its capability in multi-component simulation through thermal and mechanical analysis of prismatic fuel elements.
@@
Concretely, our contributions are threefold: (1) We introduce a novel approach, M2PDE, for multiphysics and multi-component PDE simulations, framing the problem in terms of joint probabilistic modeling. By training on decoupled (small structure) training data, it can generate coupled (large structure) solutions. (2) We create and open-source benchmark datasets for both multiphysics and multi-component PDE simulations, providing a valuable resource for future research. (3) Our method demonstrates success in both domains. For multiphysics simulation, M2PDE accurately predicts coupled solutions in complex problems where surrogate models fail. In multi-component simulations, M2PDE, trained on single components, accurately predicts larger structures with up to 64 components.
@@ 2 Related Work
Multi-component simulation. To our knowledge, there do not exist utilized machine learning methods specifically designed for multi-component simulation. A relevant study is the CoAE-MLSim algorithm (Ranade et al., 2021). This algorithm combines neural networks with numerical iteration. It first partitions the computational domain into multiple subdomains, and then trains a neural network to learn the flux conversation between subdomains, similar to the domain decomposition algorithm (Chan & Mathew, 1994). During inference, the neural network with flux conservation is applied sequentially in each subdomain, looping until convergence. We further extend this algorithm to multi-component simulation and use it as a baseline. Besides, graph neural network (GNN) (Wu et al., 2020; Fan et al., 2024) can learn on small graphs and inference on larger graphs (Xu et al., 2019); Graph Transformer (Kreuzer et al., 2021) employs the Laplacian matrix of a graph to characterize its structure, and by leveraging the Transformer architecture, it achieves learning on graphs. We also compare M2PDE with GNN and Graph Transformer.
Compositional models. Recent research has extensively explored the compositional combination of generative models for various applications, including image or video synthesis (Du et al., 2020, 2023; Po & Wetzstein, 2024; Yang et al., 2023), multimodal perception (Li et al., 2022), trajectory planning (Du et al., 2019; Urain et al., 2023), inverse design (Wu et al., 2024b), and data assimilation (Rozet & Louppe, 2023). A particularly effective approach for combining predictive distributions from local experts is the product of experts framework (Hinton, 2002; Cohen et al., 2020). Their focus is on how a single object is influenced by multiple factors, such as generating images that meet various requirements in image generation (Du et al., 2023) or enhancing the lift-to-drag ratio under the influence of two wings in inverse design (Wu et al., 2024b). However, our problem involves multiple objects, such as multiple physical processes and components, requiring the capture of interactions between these fields or components. Existing research is not applicable to multiphysics and multi-component PDE simulation. To the best of our knowledge, we are the first to introduce a compositional generative approach to multiphysics and multi-component PDE simulations, demonstrating how this framework enables generalization to far more complex simulation tasks than those encountered during training.
@@ 3 Method
In this section, we introduce the principle of M2PDE solving multiphysics and multi-component PDE simulation in section 3.1 and section 3.2, respectively.
@@ 3.1 Multiphysics simulation
Consider a complex multiphysics simulation that consists of multiple physical processes : $z=(z_{1},z_{2},\ldots,z_{N})$, each $z_{i}$ represents the trajectory of one or more physical fields that belong to that specific $i$-th physical process. For example, the mechanics might contain both the stress and strain fields in three directions over a given time interval $[0,T]$. Each process $z_{i}$ has its own governing equation which depends on other processes , and solving equations for other processes also requires that process. Therefore, all equations must be solved simultaneously to achieve the most accurate representation of the physical system.
@@
$E(z)+\mathrm{log}Z=[E(z_{i}\mid z_{\neq i})+\mathrm{log}Z(z_{\neq i})]$ (5)
$+[E(z_{\neq i})+\mathrm{log}Z_{\neq i}]$
@@
$L_{\mathrm{MSE}}=||\epsilon-\epsilon_{\theta}(\sqrt{1-\beta_{s}}z_{i}+\sqrt{\beta}_{s}\epsilon;z_{\neq i},s)||_{2}^{2}$ (7)
where the denoising network $\epsilon_{\theta}(\cdot)$ corresponds to the gradient of the energy function $\nabla_{z}E_{\theta}(\cdot)$ (Du et al., 2023), $\epsilon\sim\mathcal{N}(0,I)$. During inference, we sample from the joint distribution $p(z_{1},z_{2},...z_{N})$ via (Ho et al., 2020):
$z_{i,s-1}=\frac{1}{\sqrt{\alpha_{s}}}\left(z_{i,s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}^{i}(z_{i,s}\mid z_{\neq i}^{e},s)\right)$ (8)
$+\sigma_{s}w,\,\,\,w\sim\mathcal{N}(0,I)$
@@
for $s=S,S-1,...1$ and $i=1,2,...n$. Here, ${z}_{i}^{e}$ represents the estimated value for the $i$th field $z_{i}$, $z^{e}_{\neq i}=(z^{e}_{1},...z^{e}_{i-1},z^{e}_{i-1},...z^{e}_{n})$, and $\sigma_{s}$ is the noise level, the superscript $e$ represents estimation.
@@
Algorithm 1 Algorithm for multiphysics simulation by M2PDE.
@@
1:  $z_{i}^{e}\sim\mathcal{N}(0,I)$  // initialize estimated fields $z_{i}^{e}$// An external loop to improve the estimated fields $z_{i}^{e}$:
@@
3:      // initialize each physical fields $z_{i}$, for $i$ in $1,...,N$
4:     $\hat{z}_{i}^{e}\leftarrow z_{i}^{e}$  // previous estimated fields $\hat{z}_{i}^{e}$
5:     $z_{i}^{e}\sim\mathcal{N}(0,\bf I)$  // current estimated fields $z_{i}^{e}$
6:     $z_{i,S}\sim\mathcal{N}(0,\bf I)$  // intial physical fields $z_{i}$// denoising cycle of diffusion model:
7:     for $s=S,...,1$ do
8:        $\lambda=1-\frac{s}{S}$ if $k>1$ else 1 // define the weights of $\hat{z}_{i}^{e}$ and $z_{i}^{e}$// loops for each physical process:
9:        for $i=1,...,N$ do
10:           $w\sim\mathcal{N}(0,\bf I)$// use weighted estimated fields as conditions for single step denoising:
11:           $z_{i,s-1}=\frac{1}{\sqrt{\alpha_{s}}}(z_{i,s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}^{i}(z_{i,s}\mid\lambda z_{\neq i}^{e}+(1-\lambda)\hat{z}_{\neq i}^{e},C,s))+\sigma_{s}w$// update the estimation of current field:
12:           $z_{i}^{e}=\frac{1}{\sqrt{\overline{\alpha}_{s}}}(z_{i,s}-\sqrt{1-\overline{\alpha}_{s}}\epsilon_{\theta}^{i}(z_{i,s}\mid\lambda z_{\neq i}^{e}+(1-\lambda)\hat{z}_{\neq i}^{e},C,s))$